In an impulse turbine of the Pelton type, the jet is turned by the bucket by...

  1. In an impulse turbine of the Pelton type, the jet is turned by the bucket by 165°. The head available at the nozzle is 750 m and blade speed ratio is 0.46. Cv = 0.98. Relative velocity is reduced by 12% due to friction. Determine the hydraulic efficiency. If the flow available is 20 m3/s. What is the power potential. Assuming 5 units of equal power, determine the jet diameter and wheel diameter if d/D = 12.
Answer #1

o type - Impulse – Pellon type. ☺_Headchyle 750 m _ _blade ratio (0)= 0.16 © Co-efficient of Velocity (cv) = 0.98 ☺ bucket dedue to gravity O firstly, we calculate. Velocity of_Jot () - Vi = Cragh ::g=9.81 - = 0.98 V 269.81*750_ acceleration V = 118.due to gravity O firstly, we calculate. Velocity of_Jot () - Vi = Cragh ::g=9.81 - = 0.98 V 269.81*750_ acceleration V = 118.80 Reduchon relative velocty (uri) at Inlet due to poichon Vr (friction) = 88 % vr 124. friction 1 5.88Vors 3,88 x 63.074 m/sin = 13,022.38 --(H18-87412 n = 13,022.38 14731-02 hh - 0.9215 (24) Inn = 92.15% Answor. Pelton wheel hydraulic efficiany = 9power = 130, 223,808 4000 = 130,223.80 kw Lanswer. P Jet dia meher given jet ra discharge Q = Area ok jer x Velocity or jet .